# Active Data Curation Effectively Distills Large-Scale Multimodal Models

### v1

@@ 1 Introduction
Deploying multimodal foundation models [14] like CLIP [119] on edge devices is challenging due to their high inference costs and memory footprints. This motivates the need for compressing these foundation models into smaller ones that can be efficiently deployed for cheaper and faster inference, while retaining the performance of their larger counterparts. Knowledge distillation (KD) [65] is a classic model compression technique—a method for transferring knowledge from a large-scale “teacher” model into a smaller “student” model, via matching the student and teacher logits, features or activations. KD has been extensively deployed for creating small, performant models like Gemma-2 [148], Phi-3 [4], Gemini-1.5 Flash [126], and SD3-Turbo [136].
@@
Here, our primary goal is to downscale contrastive vision-language models such as CLIP [119] and SigLIP [190]), without compromising downstream performance. Prior works in this domain focus on complex KD strategies as the key solution—the current SoTA (TinyCLIP [173] and MobileCLIP [155]) use extensive combinations of methods such as strong data-augmentation policies, multi-teacher ensembles, synthetic captions, complicated weight-inheritance, pruning strategies, and bespoke model architectures.
In this work, we take an alternative simplifying approach—we propose to use active data curation as an effective strategy for distilling large vision-language models (VLMs) into smaller and more FLOP-efficient multimodal models.
Our method, ACID (Active Curation as Implicit Distillation), automatically selects samples that reduce the performance gap between a small model being trained (student) and a larger frozen model (reference). Under appropriate conditions, we find this to be an effective alternative to standard KD for distilling knowledge to small models. To the best of our knowledge, this is a novel and surprising finding since prior work in data curation assumed that larger models could not be used to select data for smaller ones as the capacity gap between the two could prove inhibitory [107, 48]. Through a novel theoretical interpretation and extensive experiments, instead, we demonstrate that ACID is not only effective but actually improves over KD, exhibiting more favourable scaling with respect to training compute. We also conduct careful ablation studies uncovering critical factors influencing the quality of the trained student model, including, reference model capacity and training dataset.
@@
Finally, we incorporate our findings to develop a simple yet extremely effective pretraining recipe, ACED (ACID with Explicit Distillation), to train FLOP-efficient image-text contrastive models. Our method, absent bespoke components such as efficient architectures or data-augmentations, outperforms SoTA CLIP and SigLIP models with greater FLOP-efficiency at inference-time and shows a significant improvement over 27 downstream tasks against the previous SoTA for FLOP-efficient models [155, 173]. We further demonstrate that our ACED vision-encoders provide strong backbones for generative multimodal models, outperforming larger and FLOP-inefficient vision-encoders on image-captioning and visual-question-answering (VQA) tasks.
@@ 2 Related Work
Multimodal Data Curation. Recent works have emphasised the importance of data quality for pretraining multimodal models [109, 48, 153, 41, 106]. Specifically, it has been shown that offline curation of noisy web-scale data can result in large pretraining efficiency gains [144, 69, 20, 160, 1, 19, 103, 178, 2, 101, 169, 42, 168, 138, 75, 186, 167, 159].
@@
However, such static offline curation methods that pre-filter data do not take into account the training dynamics of the current learner model, and hence can suffer at larger scales [53]. As a result, there have been many recent attempts to introduce online batch selection criteria that account for the current state of the learner (e.g., at each step select training samples that have the largest learner loss) [100, 67, 73, 143, 70, 170, 197, 137, 177]. Going further, Mindermann et al. [107] introduced the RHO-Loss that considers both current learner state and a pretrained data-selector (reference) model. This criterion has since been used in many efforts to improve the efficiency of foundation model pretraining [34, 35, 17, 66, 39, 31]. As these methods seek to improve pretraining efficiency, the pretrained reference models that are used as data selectors are typically smaller than the learner models they are used to train [34, 35, 42]. In fact, Fang et al. [42], Yu et al. [186], Gadre et al. [48] all showed that increasing reference model size can potentially hurt learner model performance. Our work tells a different story, finding that large data selectors can effectively curate data for inference-time FLOP-efficient learner models.
@@
Knowledge Distillation. First introduced by Buciluǎ et al. [18] and further popularized by Hinton [65], Ba and Caruana [7], knowledge distillation (KD) is a classic technique for transferring knowledge from a larger model (teacher) to another smaller one (student), by optimizing the student to match certain outputs (logits, features, intermediate activations etc.) of the teacher model. It has been extensively used for compressing large models into smaller, deployable ones in unimodal tasks like image-classification [12, 25, 158, 165, 113, 149] and language representation learning [179, 5, 134, 76, 95, 55, 146]. Further works have extended KD to use teacher-ensembles [141, 21, 185, 145, 105, 37, 200, 135], and different distillation training objectives [68, 196, 175, 147, 151, 122, 92].
@@
Most relevant to our work are KD methods in the mutlimodal regime, which is an underexplored area. Nevertheless, there are a number of recent efforts to distill strong but efficient CLIP models.  Sameni et al. [133] introduced SF-CLIP, a method using masked distillation, while Vasu et al. [155] proposed MobileCLIP, exploring downscaling CLIP models for mobile-deployment by using multi-teacher contrastive-KD, synthetic captions, and data-augmentations. Wu et al. [173] further proposed TinyCLIP—a weight inheritance method combined with an affinity-mimicking strategy to yield tiny CLIP models. Yang et al. [180] conducted an empirical study (CLIP-KD) on different objective functions for effectively distilling CLIP models, across different scales. Finally, CLIP-CID [183] uses an image semantic balancing strategy coupled with cluster-instance discrimination for better teacher-to-student knowledge transfer during the KD process. We compare against all of these methods in our experimental results in Sec. 4.
@@
Accelerating Knowledge Distillation There has been some prior work attempting to make KD-based pretraining more efficient [142, 140, 188]. Some works have investigated accelerating vanilla KD using active learning in small-scale classification tasks [163, 83, 176]. However, these approaches require a costly iterative process, involving synthetic generation, followed by active sample selection to produce pseudo-labels from a teacher model, thereby limiting their scalability. Another line of work has studied data-selection methods for improving KD, typically using uncertainty-based data, logit and feature selection [123, 130, 161, 172, 97, 199, 90, 59, 162], contextual retrieval and sample augmentation from a large data pool [98, 71, 94, 50, 191, 118], or influence-function based sample selection [83, 184]. Contrary to these works, Beyer et al. [12] and Hao et al. [57] suggest that vanilla knowledge distillation is optimal in “infinite-data regimes”.
@@
Surprisingly, all these prior works operate primarily in the unimodal image/text classification regime, and none have been scaled to multimodal foundation model training. We showcase, for the first time, that simple data selection using online batch selection outperforms standard KD for pretraining multimodal models. We further study the optimal strategies for combining vanilla KD and active data curation in order to best leverage their complementary strengths.
@@ 3.2 ACID: Active Curation as Implicit Distillation
Active Data Curation as Implicit Distillation (ACID). We now show formally that active curation can be cast as “implicit distillation” and should benefit from larger reference models. The model now minimizes ^$\hat{\mathcal{L}}=\frac{1}{b}\sum_{x_{i}\sim\mathcal{A}[\mathcal{S}]}\mathcal{L}(x_{i};\mathcal{B})$, which in expectation is $\mathcal{E}=\mathbb{E}[\hat{\mathcal{L}}]=\sum_{x\in\mathcal{D}}a(x)\mathcal{L}(x;\mathcal{B})$ given that super-batches $\mathcal{S}$ are sampled uniformly. Recall that $\mathcal{L}(x;\mathcal{B})=-\sum_{i=1}^{b}y_{i}(x)\log q_{i}(x)$, where $y_{i}$ are the labels of the contrastive task and $q_{i}$ are the probabilities induced by the pairwise similarities of the student $\theta$. Let $p_{i}$ be the probabilities induced by the reference model $\theta_{\text{ref}}$. In the case of easy-reference scoring and the softmax loss, $a(x)=\frac{1}{Z}\exp\sum_{i=1}^{b}y_{i}(x)\log p_{i}(x)=\frac{1}{Z}p_{i^{*}}(x)$ where $i^{*}$ is the index of the one-hot label $y(x)$. We derive the following equality (see supplementary for details),
@@
i.e. implicit distillation on hard examples (“H-ACID”) according to the student (see supplementary for details). Prioritizing high-loss examples has been shown to reliably accelerate learning in settings where targets are high-quality [100], as is the case with the combined targets in our ACID.
@@
Table 1: Method Instantiations. We recover different method configurations from our unified objective (see Sec. 3.3), by specifying the data-selection strategies across different batches as well as hyperparameter values. We further indicate the effective mini-batch size per-iteration used by each method for training, and colour-code the different methods for easy referencing from Sec. 4.
@@ 4.1 Implementation Details
Model Architecture and Sizes. Unless otherwise specified, we use standard ViT-S [33] and BERT-small [32] models as our student image-text encoders (typically used by open_clip). For some student ablations, we also use (ViT-Ti image, Ti text) and (ViT-B image, B text) configurations. For our reference and teacher models, we conduct a sweep over different sizes—(ViT-Ti, Ti), (ViT-S, S), (ViT-B, B), (ViT-L, L), (ViT-H, H), and (ViT-g, g) for (image, text) encoders respectively. We pretrain all our models ($\theta_{\text{teacher}}$, $\theta_{\text{ref}}$, $\theta$) from scratch. For more details, refer to supplementary.
@@
Evaluation Protocol: StableEval. We evaluate our models on a diverse set of benchmarks including zero-shot classification and image-text retrieval datasets following prior multimodal pretraining work [48, 84, 178]. However, many works select non-standardized sets of evaluations and fail to sufficiently justify the reliability of the evaluations they use. To rigorously define an evaluation suite, we collate a standard list of 34 candidate evaluations and conduct a systematic analysis of their reliability. By repeating the same canonical pretraining run multiple times (e.g., CLIP pretraining on DataComp with the exact same data ordering, see supplementary for details), we evaluate the variability of each metric across random seeds. In Fig. 3 (left), we find an extreme range in variability across evaluations (stds from 0.15% to 12.5%) which hinders comparisons among different methods. Inspired loosely by the continuous inverse-variance weighting (IVW) method for minimizing variance of aggregated random variables [58], we develop a method for choosing a discrete, stable subset of relevant evaluations. We compute the variability of a progressively growing set of evaluations, starting from least variable and incrementally adding more variable ones, in ascending order. For a subset of size $N$, ${std}{(E_{1}...E_{N})}{=}{\sqrt{\frac{1}{N^{2}}\sum_{i}{var(E_{i})}}}$. Because of the $1/N^{2}$ scaling, adding more datasets decreases the variability of the average (Fig. 3 (right)) to a critical point. However, adding highly variable evaluations outweighs this term, increasing the average variability. We limit the evaluation set to remain highly reliable (i.e. with lower variability than the most reliable individual evaluation (${<}{0.15}$)) while still including as many evaluations possible to maximize coverage and diversity, yielding the 27 StableEval set.
@@
Training Configurations. We implement our model with the big_vision codebase [11], following the contrastive pretraining setup of SigLIP [190]. Unless otherwise specified, we train for $3$ billion total samples seen, with a batch-size of ${b}{=}{32,678}$ with the sigmoid contrastive loss (Eq. 7). The image-encoder takes images resized to $({256}{\times}{256})$ without any additional augmentations. The text-encoder uses a sentencepiece tokenizer [80] trained on English-C4 [120], with a vocabulary size of $32,000$. We truncate all text captions to the first $64$ tokens. For most experiments, we use an rsqrt learning rate scheduler [189], with a peak learning-rate of $0.001$, and linear-warmup and linear-cooldown applied for $10\%$ of total steps. By default, we use a filtering ratio of ${f}{=}{0.8}$ when using ACID sampling, leading to a super-batch-size of ${B}{=}{163,840}$. We sweep over ${\lambda}{=}{\{{0.5}{,}{1.0}{,}{2.0}\}}$ for finding the optimal loss-weight for the Softmax-KD loss (Eq. 3). For more details, refer to supplementary.
@@ 4.2.2 ACID outperforms standard distillation
Having demonstrated the scaling behaviour of ACID and showcasing favourable performance comparisons to Softmax-KD using a single reference/teacher model dataset, we next demonstrate that ACID outperforms explicit distillation methods across different teacher/reference pretraining datasets, distillation objectives, and student model sizes.
@@
Different Distillation Objectives. Prior work has explored several different objectives for multimodal distillation, beyond standard Softmax-KD. Here, we compare our ACID method to some of these, including a Sigmoid-KD loss [173] and a Feature-Matching KD loss [180] (see supplementary for details). Further, the SoTA multimodal distillation method, CLIP-KD [180], advocates combining these losses for best performance. We therefore also compare against two combination methods—Softmax+Sigmoid and Softmax+Feature-Matching. In Fig. 5 (center), we show that ACID, without any additional complexity, still comprehensively outperforms all of the other distillation objectives.
@@ 4.3 ACED: ACID and KD are complementary
ACED—How to combine? We now discuss strategies for combining ACID and Softmax-KD. The simple strategy, ACIDistill, samples a training batch using ACID and applies both the contrastive and softmax-distillation loss on that batch. An alternative strategy, IIDistill, samples two batches independently, one with ACID sampling and the other IID sampled, and applies the distillation loss on the IID batch while the contrastive loss is applied on the ACID batch. We study the scaling behaviour of both these strategies by training a ViT-S student with a WebLI-L teacher and a WebLI-curated++-reference model, for 3B, 6.5B and 13B samples seen. We observe that the ACIDistill method showcases better performance across all compute budget scales (see supplementary). Hence, going forward, we use ACIDistill as the default strategy for combining ACID and Softmax-KD, and refer to that as our main ACED method.
@@
Table 2: ACED outperforms all prior state-of-the-art methods. We showcase results for our method at three different model inference-GFlop scales. Across all three model scales, our ACED models’ performance improves on the prior SoTA across the 27 StableEval evaluations, while using fewer inference FLOPs. For better interpretability, we also break down the evaluations into individual benchmarks (e.g., ImageNet, COCO) and groupings (e.g., ImageNet distribution shifts, other object categorization, and scene classification—for details, see supplementary). MobileCLIP samples seen include two captions for each image, effectively doubling the total unique pairs.∗∗TinyCLIP models are not trained from scratch, but use a complex weight inheritance strategy from pretrained models.
@@
Table 3: LiT-Decoder Evaluations.

### v2

@@ 1 Introduction
Deploying multimodal foundation models [14] like CLIP [119] on edge devices is challenging due to their high inference costs and memory footprints. This motivates the need for smaller, inference-efficient models that retain the performance of their larger counterparts. Knowledge distillation (KD) [65] is a classic model compression technique—a method for transferring knowledge from a large-scale “teacher” model into a smaller “student” model, via matching student and teacher logits, features or activations. KD has been extensively deployed for creating small, performant models like Gemma-2 [148], Phi-3 [4], Gemini-1.5 Flash [126], and SD3-Turbo [136].
@@
Here, our primary goal is to downscale contrastive vision-language models , without compromising downstream performance. Prior works in this domain focus on complex KD strategies as the key solution—the current SoTA (TinyCLIP [173] and MobileCLIP [155]) use combinations of methods such as strong data-augmentation policies, multi-teacher ensembles, synthetic captions, weight-inheritance, weight-pruning, and bespoke model architectures.
In this work, we seek a simplified approach. Specifically, we propose using active data curation as an effective strategy for distilling large vision-language models (VLMs) into smaller and more FLOP-efficient multimodal models.
Our method, ACID (Active Curation as Implicit Distillation), automatically selects samples that reduce the performance gap between a small student model and a larger reference model. Under appropriate conditions, we find this is a surprisingly effective distillation approach. To the best of our knowledge, this is a novel finding since prior works in data curation assume that larger models can not be used to select data for smaller ones, due to the capacity gap [107, 48]. Through a novel theoretical interpretation and extensive experiments, instead, we demonstrate that ACID is not only effective but also improves over standard KD, exhibiting more favourable scaling with respect to training compute. We also conduct careful ablation studies that uncover factors influencing the quality of the trained student model, including, reference model capacity and training dataset.
@@
Based on this finding, we propose our final pretraining recipe, ACED (ACID with Explicit Distillation), andtrain very strong FLOP-efficient image-text contrastive models. Our method, absent bespoke components such as efficient architectures or data-augmentations, outperforms SoTA CLIP and SigLIP models with greater FLOP-efficiency at inference-time and shows a significant improvement over 27 downstream tasks against these prior SoTA FLOP-efficient models [155, 173]. We further demonstrate that our ACED vision-encoders provide strong backbones for generative multimodal models, outperforming larger and FLOP-inefficient vision-encoders on image-captioning and visual-question-answering (VQA) tasks.
@@ 2 Related Work
Multimodal Data Curation. Recent works have emphasised the importance of data quality for multimodal pretraining [109, 48, 153, 41, 106]. Specifically, offline curation of noisy web-scale data can result in large pretraining efficiency gains [69, 20, 160, 1, 19, 103, 178, 2, 101, 169, 42]. However, these static methods that pre-filter data do not take into account the training dynamics of the current learner model. As a result, there have been many recent attempts to introduce online batch selection criteria that account for the current state of the learner (e.g., at each step select training samples that have the largest learner loss) [100, 67, 73, 143, 70, 170, 197, 137, 177]. The RHO-Loss [107] goes further to consider current learner state and a pretrained data-selector (reference) model. This criterion has since been used in many efforts to improve the efficiency of foundation model pretraining [34, 35, 17, 66, 39, 31]. As these methods seek to improve pretraining efficiency, the pretrained reference models that are used as data selectors are typically smaller than the learner models they are used to train [34, 35, 42]. In fact, many works have shown that increasing reference model size can potentially hurt learner model performance [42, 186, 48] . Our work tells a different story, finding that large data selectors can effectively curate data for inference-time FLOP-efficient learner models.
@@
Knowledge Distillation. First introduced by Buciluǎ et al. [18] and further popularized by Hinton [65], Ba and Caruana [7], knowledge distillation (KD) is a classic technique for transferring knowledge from a larger model (teacher) to a smaller one (student) by optimizing the student to match outputs of the teacher. Such methods have been used for compressing large models in unimodal tasks like image-classification [12, 25, 158, 165, 113, 149] and language representation learning [179, 5, 134, 76, 95, 55, 146]. Further works have extended KD to use teacher-ensembles [141, 21, 185, 145, 105, 37, 200, 135], and different distillation training objectives [68, 196, 175, 147, 151, 122, 92].
@@
Most relevant to our work, there are a number of recent efforts to distill CLIP models. SF-CLIP [133] explores masked distillation, while MobileCLIP [155] uses multi-teacher contrastive-KD, synthetic captions, and data-augmentations. TinyCLIP [173] proposes a weight inheritance method combined with an affinity-mimicking strategy. An empirical study (CLIP-KD [180]) also has explored different objective functions for effectively distilling CLIP models, across different scales. Finally, CLIP-CID [183] uses an image semantic balancing strategy coupled with cluster-instance discrimination for better teacher-to-student knowledge transfer during the KD process. We compare against all of these methods in our experimental results in Sec. 4.
@@
Accelerating Knowledge Distillation. Prior works have investigated accelerating vanilla KD using active learning in small-scale settings [163, 83, 176]. However, these approaches require a costly iterative process, involving synthetic generation, followed by active sample selection to produce pseudo-labels from a teacher model, thereby limiting their scalability. Other works have studied data-selection methods for improving KD, typically using uncertainty-based data, logit and feature selection [123, 130, 161, 172, 97, 199, 90, 59, 162], contextual retrieval and sample augmentation from a large data pool [98, 71, 94, 50, 191, 118], or influence-function based sample selection [83, 184]. Contrary to these works, others suggest that vanilla knowledge distillation is optimal in “infinite-data regimes” [12, 57]. Surprisingly, these studies operate primarily in the unimodal image/text classification regime, and none have been scaled to multimodal foundation model training.
@@
We showcase, for the first time, that simple data selection using online batch selection outperforms standard KD for pretraining multimodal models. We further study the optimal strategies for combining vanilla KD and active data curation in order to best leverage their complementary strengths.
@@ 3.2 ACID: Active Curation as Implicit Distillation
Active Data Curation as Implicit Distillation (ACID). We now show formally that active curation can be cast as “implicit distillation” and should benefit from larger reference models. The model now minimizes ^$\hat{\mathcal{L}}=\frac{1}{b}\sum_{x_{i}\sim\mathcal{A}[\mathcal{S}]}\mathcal{L}(x_{i};\mathcal{B})$, which in expectation is $\mathcal{E}=\mathbb{E}[\hat{\mathcal{L}}]=\sum_{x\in\mathcal{D}}a(x)\mathcal{L}(x;\mathcal{B})$ given that super-batches $\mathcal{S}$ are sampled uniformly. Recall that $\mathcal{L}(x;\mathcal{B})=-\sum_{i=1}^{b}y_{i}(x)\log q_{i}(x)$, where $y_{i}$ are the labels of the contrastive task and $q_{i}$ are the probabilities induced by the pairwise similarities of the student $\theta$. Let $p_{i}$ be the probabilities induced by the reference model $\theta_{\text{ref}}$. In the case of easy-reference scoring and the softmax loss, $a(x)=\frac{1}{Z}\exp\sum_{i=1}^{b}y_{i}(x)\log p_{i}(x)=\frac{1}{Z}p_{i^{*}}(x)$ where $i^{*}$ is the index of the one-hot label $y(x)$. We derive the following equality (see Appendix C for details),
@@
i.e. implicit distillation on hard examples (“H-ACID”) according to the student (see Appendix C for details). Prioritizing high-loss examples has been shown to reliably accelerate learning in settings where targets are high-quality [100], as is the case with the combined targets in our ACID.
@@
Table 1: Method Instantiations recovered from our unified objective (see Sec. 3.3), by specifying data-selection strategies across different batches and hyperparameter values. We further indicate the effective mini-batch size per-iteration used by each method, and colour-code different methods for easy referencing from Sec. 4.
@@ 4.1 Implementation Details
Model Architecture and Sizes. Unless otherwise specified, we use standard ViT-S [33] and BERT-small [32] models as our student image-text encoders. For some student ablations, we also use (ViT-Ti image, Ti text) and (ViT-B image, B text) configurations. For our references and teachers, we sweep over different sizes—(ViT-Ti, Ti), (ViT-S, S), (ViT-B, B), (ViT-L, L), (ViT-H, H), and (ViT-g, g) for (image, text) encoders respectively. We pretrain all our models ($\theta_{\text{teacher}}$, $\theta_{\text{ref}}$, $\theta$) from scratch. For more details, refer to Appendix F.
@@
Evaluation Protocol: StableEval. We evaluate our models on a diverse set of benchmarks including zero-shot classification and image-text retrieval datasets following prior multimodal pretraining works [48, 84, 178]. However, many works select non-standardized sets of evaluations and fail to sufficiently justify the reliability of the evaluations they use. To rigorously define an evaluation suite, we collate a standard list of 34 candidate evaluations and conduct a systematic analysis of their reliability. By repeating the same canonical pretraining run multiple times (e.g., CLIP pretraining on DataComp with the exact same data ordering, see Appendix A for details), we evaluate the variability of each metric across random seeds. In Fig. 3 (left), we find an extreme range in variability across evaluations (stds from 0.15% to 12.5%) which hinders comparisons among different methods. Inspired loosely by the continuous inverse-variance weighting (IVW) method for minimizing variance of aggregated random variables [58], we develop a method for choosing a discrete, stable subset of relevant evaluations. We compute the variability of a progressively growing set of evaluations, starting from least variable and incrementally adding more variable ones, in ascending order. For a subset of size $N$, ${std}{(E_{1}...E_{N})}{=}{\sqrt{\frac{1}{N^{2}}\sum_{i}{var(E_{i})}}}$. Because of the $1/N^{2}$ scaling, adding more datasets decreases the variability of the average (Fig. 3 (right)) to a critical point. However, adding highly variable evaluations outweighs this term, increasing the average variability. We limit the evaluation set to remain highly reliable (i.e. with lower variability than the most reliable individual evaluation (${<}{0.15}$)) while still including as many evaluations possible to maximize coverage and diversity, yielding the 27 StableEval set.
@@
Training Configurations. Unless otherwise specified, we train for $3$ billion total samples seen, with a batch-size of ${b}{=}{32,678}$ with the sigmoid contrastive loss (Eq. 7). The image-encoder takes images resized to $({256}{\times}{256})$ without additional augmentations. The text-encoder uses a sentencepiece tokenizer [80] trained on English-C4 [120], with a vocabulary size of $32,000$. We truncate all text captions to the first $64$ tokens. For most experiments, we use an rsqrt learning rate scheduler [189], with a peak learning-rate of $0.001$, and linear-warmup and linear-cooldown applied for $10\%$ of total steps. By default, we use a filtering ratio of ${f}{=}{0.8}$ when using ACID sampling, leading to a super-batch-size of ${B}{=}{163,840}$. We sweep over ${\lambda}{=}{\{{0.5}{,}{1.0}{,}{2.0}\}}$ for finding the optimal loss-weight for the Softmax-KD loss (Eq. 3). For more details, refer to Appendix E.
@@ 4.2.2 ACID outperforms standard distillation
Having demonstrated the favourable scaling behaviour of ACID vs. Softmax-KD using a single teacher/reference model dataset, we next demonstrate that ACID outperforms explicit distillation across different teacher/reference pretraining datasets, objective functions, and student model sizes.
@@
Different Distillation Objectives. Prior works have explored several different objectives for multimodal distillation, beyond standard Softmax-KD. Here, we compare our ACID method to some of these, including a Sigmoid-KD loss [173] and a Feature-Matching KD loss [180] (see Appendix D for more details). Further, the SoTA multimodal distillation method, CLIP-KD [180], advocates combining these losses for best performance. We therefore also compare against two combination methods—Softmax+Sigmoid and Softmax+Feature-Matching. In Fig. 5 (center), we show that ACID, without any additional complexity, still comprehensively outperforms all of the other distillation objectives.
@@ 4.3 ACED: ACID and KD are complementary
ACED—How to combine? We now discuss strategies for combining ACID and Softmax-KD. The simple strategy, ACIDistill, samples a training batch using ACID and applies both the contrastive and softmax-distillation loss on that batch. The alternative, IIDistill, samples two batches independently, one with ACID sampling and the other IID sampled, and applies the distillation loss on the IID batch while the contrastive loss is applied on the ACID batch. We study the scaling behaviour of both strategies by training ViT-S students with WebLI-L teachers and WebLI-curated++-references, for 3B, 6.5B and 13B samples seen. We observe ACIDistill showcases better performance across all compute budget scales (see Sec. H.1). Hence, going forward, we use ACIDistill as the default strategy for combining ACID and Softmax-KD, and refer to that as our main ACED method.
@@
Table 2: ACED outperforms all prior state-of-the-art methods. We showcase results for our method at three different model inference-GFlop scales. Across all three model scales, the performance of our ACED models improves on the prior SoTA across the 27 StableEval evaluation datasets, while using fewer inference FLOPs. For better interpretability, we also break down the evaluations into individual benchmarks (e.g., ImageNet, COCO) and groupings (e.g., ImageNet distribution shifts, other object categorization, and scene classification—for details, see Appendix A). MobileCLIP samples seen include two captions for each image, effectively doubling the total unique pairs.∗∗TinyCLIP models are not trained from scratch, but use a complex weight inheritance strategy from pretrained models.
@@
Table 3: LiT-Decoder Evaluations. Our ACED vision-encoders also improve performance of multimodal-decders on captioning and VQA tasks, when compared to strong SigLIP and SiLC baselines.